OpenStudy (swaqqout_kid1):

The product of an initial investment, I, and the quantity of the sum of 1 and the annual interest rate, r, raised to the power of n, the number of years of the investment, is equal to M, the current amount of money in an investment account. If an initial investment of $1,028.00 is made to an account with an annual interest rate of 9%, what will be the value of M after 4 years? Round to the nearest cent. @shinalcantara

OpenStudy (shinalcantara):

According to the first paragraph, the formula would be \[M = I(1+r)^n\] Substitute the given values. Can you do that @SwaqqOut_KID1

OpenStudy (shinalcantara):

I = 1,028 r = 9% or in decimal 0.09 n = 4 ------------- Substitute that and you'll get M

OpenStudy (swaqqout_kid1):

wat am i suppose to do again

OpenStudy (shinalcantara):

\[M = I(1+r)^n\] I = 1,028 r = 9% or in decimal 0.09 n = 4 ---------------- Substitute the values and you'll get M --------------- \[M = 1,028(1+0.09)^4\] Solve for M